What is the next three terms of 4,8,16,32,64

Answer by greenestamps(13200)

There is no way to know what the next term(s) are; ANY next three numbers will make a valid sequence.

You can look for an APPARENT pattern in the numbers to GUESS a POSSIBLE or even PROBABLE answer; but there is no way to know whether any answer you come up with is correct.

In this example, each term is double the previous term, SUGGESTING a POSSIBLE answer:

4*2 = 8
8*2 = 16
16*2 = 32
32*2 = 64
64*2 = 128
128*2 = 256
256*2 = 512

But, again.... Despite how OBVIOUS that answer is, it is only a GUESS.